Nguyen Inc. is considering the purchase of a new computer system (ICX) for $180,000. The system will require an additional $40,000 for installation. If the new computer is purchased it will replace an old system that has been fully depreciated. The new system will be depreciated over a period of 8 years using straight-line depreciation. If the ICX is purchased, the old system will be sold for $10,000. The ICX system, which has a useful life of 8 years, is expected to increase revenues by $32,000 per year over its useful life. Operating costs are expected to decrease by $5,000 per year over the life of the system. The firm is taxed at a 40 percent marginal rate. Round your answers to the nearest dollar.

What net investment is required to acquire the ICX system and replace the old system? $

Compute the annual net cash flows associated with the purchase of the ICX system. $
Answer #1

1) Net investment = Purchase price of new computer system + Installation charges - cash flow from sale of old system
Purcashe price of new computer system = $ 180000
Installation charges = $ 40000
Cash flow from sale of old asset = sales price (1-tax rate)
= 10,000(1-40%)
= 10000(1-0.4)
= 10000(0.6)
= 6000 $
Net Investment = 180000+40000-6000
=214000 $

2) Statement showing annual cash flow

Particulars 1 2 3 4 5 6 7 8
Increase in revenue 32000 32000 32000 32000 32000 32000 32000 32000
Decrease in operating cost 5000 5000 5000 5000 5000 5000 5000 5000
Depreciation
(180000+40000)/8
-27500 -27500 -27500 -27500 -27500 -27500 -27500 -27500
PBT 9500 9500 9500 9500 9500 9500 9500 9500
Tax @ 40% -3800 -3800 -3800 -3800 -3800 -3800 -3800 -3800
PAT 5700 5700 5700 5700 5700 5700 5700 5700
Add: Depreciation 27500 27500 27500 27500 27500 27500 27500 27500
Annual cash flow 33200 33200 33200 33200 33200 33200 33200 33200

Thus Annual cash flow = 33200$